# A Dual-Perspective NLG Meta-Evaluation Framework with Automatic Benchmark and Better Interpretability

### v1

@@ 5.1 Benchmarks
5.2 LLMs for LLM-as-a-Judge
We select a wide range of representative general-purpose LLMs, as well as those fine-tuned specifically for evaluation, for the LLM-as-a-judge method in our experiments. The general-purpose LLMs include GPT-4o, GPT-4o mini, GPT-4 Turbo, GPT-3.5 Turbo (Achiam et al., 2023), DeepSeek-V3 (Liu et al., 2024a), Llama-3.1-70B-Instruct (Dubey et al., 2024), Qwen-2.5-72B-Instruct (Yang et al., 2024), Gemma-2-27B-Instruct (Team et al., 2024), and Phi-4-14B (Abdin et al., 2024). The specifically fine-tuned LLMs include Auto-J-13B (Li et al., ), CRITIQUELLM-6B (Ke et al., 2024), Prometheus-13B (Kim et al., 2023), Prometheus-2-7B, Prometheus-2-8x7B (Kim et al., 2024b), Themis-8B (Hu et al., 2024b) and CompassJudger-32B (Cao et al., 2024). The detailed experimental settings are described in Appendix A.4.
@@ 5.4 Local Perspective
Furthermore, we experiment with a more intuitive evaluation approach, where LLMs directly compare each target pair with $I(1)$. As shown in Table 10, it is surprising that the LLMs except those specifically fine-tuned perform worse with direct pairwise comparison than the previous evaluating each target separately and then comparing the scores, which contradicts observations from prior studies (Liusie et al., 2024; Chen et al., 2024). We suppose this is because different evaluation approaches have their respective limits on discriminative power, and differences within the target pairs with $I(1)$ are subtle. Therefore, we further explore target pairs with greater differences up to $I(5)$, as well as other scoring ranges. The results in Figure 3 show that the direct comparison approach is only competitive when differences between targets are significant. Moreover, LLMs exhibit similar performance when the scoring range extends beyond 1–10, and the optimal range varies.
@@ 6 Related Work
As LLM-as-a-judge methods become increasingly prevalent, prior studies have designed some meta-evaluation benchmarks to assess their performance. Most of these benchmarks focus on the consistency between the evaluation results from LLMs and humans. Generally, their outputs are collected from generation models on various instruction-following tasks and humans are required to evaluate the quality of these outputs on different aspects, using direct scoring (Ye et al., 2024b; Kim et al., 2024a) or pairwise comparisons (Zheng et al., 2023; Zeng et al., 2024; Zhang et al., 2023b). A slightly different approach is taken by Thakur et al. (2024), who conduct the assessment through an objective knowledge-based question-answering task. Moreover, Tan et al. (2024) propose a pipeline for automatically constructing benchmarks focused only on factual and logical correctness.
In addition to consistency with human judgments, Wang et al. (2024a) evaluate the position bias in pairwise comparisons from LLMs. Furthermore, Ye et al. (2024a) employ the perturbation method to provide a more comprehensive quantification of 12 different biases in LLM-based evaluators. Lee et al. (2025), on the other hand, examine the consistency of evaluation results provided by LLM-based evaluators under multiple samplings and different scoring scales. And Zhao et al. (2024) consider additional expectations for the pairwise evaluation, such as transitivity.
Specifically for NLG tasks, Bavaresco et al. (2024) investigate the performance of different LLM-based evaluators across 20 NLP tasks on existing benchmarks, including translation and summarization. Wang et al. (2024c) design perturbation tests for six NLG tasks to benchmark the evaluation abilities of LLMs. Besides, some studies specifically evaluate the capabilities of LLM-based evaluators on specific NLG tasks, including translation (Freitag et al., 2024), text summarization (Liu et al., 2024c; Siledar et al., 2024), dialogue generation (Zhang et al., 2023a; Mendonça et al., 2024), and story generation (Chhun et al., 2024).
@@ A.2 Ordinal Classification Metric
Given an ordinal classification task that contains $C$ ordinal classes represented by consecutive integers, let $c_{ij}$ denote the number of items whose gold class is $j$, predicted as $i$ ($i,j\in C$), and $c_{\bullet j}=\sum_{i}c_{ij}$, $c_{i\bullet}=\sum_{j}c_{ij}$. Then the linear weighted $\kappa$ is defined as:
$\kappa=1-\frac{\sum_{j\in C}\sum_{i\in C}w_{ij}c_{ij}}{\sum_{j\in C}\sum_{i\in C}w_{ij}e_{ij}}$
where $e_{ij}=\frac{c_{i\bullet}c_{\bullet j}}{\sum_{j}\sum_{i}c_{ij}}$ represents the expected agreements when the true and predicted labels are independent, and $w_{ij}=|i-j|$ represents the weight to penalize misclassification. Amigó et al. (2020) and Sakai (2021) conduct extensive analysis and find that linear weighted $\kappa$ demonstrates the overall best evaluation capability and robustness on ordinal classification, surpassing other common metrics like accuracy and correlation.
@@ A.3 Preliminary Study on SummEval
For our proposed new meta-evaluation framework, we retain only the targets where all three human annotators provide consistent ratings in the original benchmark. For global-perspective meta-evaluation, the results are calculated on an ordinal classification task with five categories and linear weighted $\kappa$. For local-perspective meta-evaluation, since we lack a fine-grained measure for quality, the target sequence is simply constructed from targets with different original ratings, which actually contradicts our original intent. The corresponding experimental settings are similar to those described in Appendix A.4.
@@ C.1 Construction Process
Since the references in many existing NLG evaluation benchmarks are rule-based rather than human-written, we use the prompts in Table 11 and GPT-4o for generating reference targets for each source to ensure the quality. For global-perspective and local-perspective meta-evaluation, we employ OpenAI o1 to conduct the error injection using the prompts in Tables 17 and 18, respectively. Each error type is randomly selected from the decomposed evaluation sub-aspects in Tables 12 and 13.
@@ C.3 Benchmark Quality
In Table 8, we present the overall comparison results between newly constructed targets and the corresponding targets in the anchor set, based on the estimation method introduced in Section 4.2. More specifically, for the new target $t$ with the rating $r$, columns of $t^{r-1}\prec t$ and $t\prec t^{r+1}$ represent the probability that $t$ has a higher quality than targets $t^{r-1}$ with the rating $r-1$ in the anchor set $A_{r-1}$ and the probability that $t$ has a lower quality than targets $t^{r+1}$ with the rating $r+1$ in the anchor set $A_{r+1}$, respectively. Both these two probabilities are ideally as high as possible. Columns of $t^{r}\prec t$ and $t\prec t^{r}$, on the other hand, indicate the probability that $t$ has a higher and lower quality than targets $t^{r}$ with the same rating $r$ in the anchor set $A_{r}$, respectively, and these two probabilities are expected to be as close as possible. The results demonstrate that our constructed global meta-evaluation benchmarks basically align with our expectations.
Aspect $t^{r-1}\prec t$ $t^{r}\prec t$ $t\prec t^{r}$ $t\prec t^{r+1}$
@@ C.4 Construction Cost
LLM Scoring Comparison
GPT-4o 0.669 0.401
GPT-4o mini 0.641 0.500
GPT-4 Turbo 0.673 0.533
GPT-3.5 Turbo 0.599 0.341
DeepSeek-V3 0.662 0.654
Llama-3.1-70B 0.602 0.499
Qwen-2.5-72B 0.657 0.553
Gemma-2-27B 0.655 0.388
Phi-4-14B 0.652 0.561
Auto-J-13B 0.544 0.563
CRITIQUELLM-6B 0.609 0.398
Prometheus-2-7B 0.546 0.567
Prometheus-2-8x7B 0.575 0.694
CompassJudger-32B 0.634 0.639
Table 10: The accuracy of different LLMs using two evaluation approaches on the local-perspective benchmark of SummEval: scoring two targets separately followed by comparing the scores and direct pairwise comparison.
@@ Appendix D Additional Experimental Results
Consistency: Measure whether the facts in the summary are consistent with the facts in the article. Consider whether the summary does reproduce all facts accurately and does not make up untrue information. Factual Accuracy: Each fact stated in the summary accurately reflects the corresponding fact from the news article, without distorted or fabricated information.
@@
Fluency: Measure the quality of individual sentences of the summary, whether they are well-written and grammatically correct. Consider the quality of individual sentences. Grammatical Correctness: The summary adheres to standard grammar rules without errors in subject-verb agreement, capital letters, tense consistency, or word order.
@@
Relevance: Measure how well the summary captures the key points of the article. Consider whether all and only the important aspects are contained in the summary. Coverage of Important Information: The summary contains all the key points and information from the news article, without omission (e.g., removing some essential details).
@@
Understandability: Is the response understandable given the previous dialogue context? (Not if it’s on topic, but for example, if it uses pronouns, they should make sense.) Logical Flow: The response is organized in a logical sequence, ensuring smooth and clear transitions between points of the response itself.
@@
Naturalness: Does the response seem to be something that a person would naturally say? Grammatical Correctness: The response adheres to standard grammar rules without errors in subject-verb agreement, capital letters, tense consistency, word order, or spellings.
@@
Context Maintenance: Does the response serve as a valid continuation of the dialogue context (conversation history)? Logical Consistency: The response should logically follow the dialogue context to maintain a smooth continuity and have no contradictions with prior statements or facts in the dialogue context.
@@
Interestingness: Is the response dull or interesting? Content Novelty: The response introduces fresh, unexpected, or unique points and perspectives, which are different from those within the dialogue context, without just repeating the content that the dialogue has mentioned.
@@
Knowledge Use: Given the fact that the response is conditioned on, how well does the response use that fact? Fact Utilization Accuracy: The response should accurately and flawlessly use the information from the given fact, and it must not contain content that conflicts with or distorts and fabricates the given fact.

### v2

@@ 5.1 Benchmarks
5.2 LLMs for LLM-as-a-Judge
We select a wide range of representative general-purpose LLMs, as well as those fine-tuned specifically for evaluation, for the LLM-as-a-judge method in our experiments. The general-purpose LLMs include GPT-4o, GPT-4o mini, GPT-4 Turbo, GPT-3.5 Turbo (Achiam et al., 2023), DeepSeek-V3 (Liu et al., 2024a), Llama-3.1-70B-Instruct (Dubey et al., 2024), Qwen-2.5-72B-Instruct (Yang et al., 2024), Gemma-2-27B-Instruct (Team et al., 2024), and Phi-4-14B (Abdin et al., 2024). The specifically fine-tuned LLMs include Auto-J-13B (Li et al., ), CRITIQUELLM-6B (Ke et al., 2024), Prometheus-13B (Kim et al., 2023), Prometheus-2-7B, Prometheus-2-8x7B (Kim et al., 2024b), Themis-8B (Hu et al., 2024b) and CompassJudger-32B (Cao et al., 2024). The detailed experimental settings, including the evaluation instructions, are described in Appendix A.4.
@@ 5.4 Local Perspective
LLM Scoring Comparison
GPT-4o 0.669 0.401
GPT-4o mini 0.641 0.500
GPT-4 Turbo 0.673 0.533
GPT-3.5 Turbo 0.599 0.341
DeepSeek-V3 0.662 0.654
Llama-3.1-70B 0.602 0.499
Qwen-2.5-72B 0.657 0.553
Gemma-2-27B 0.655 0.388
Phi-4-14B 0.652 0.561
Auto-J-13B 0.544 0.563
CRITIQUELLM-6B 0.609 0.398
Prometheus-2-7B 0.546 0.567
Prometheus-2-8x7B 0.575 0.694
CompassJudger-32B 0.634 0.639
Table 5: The results of different LLMs using two evaluation approaches on the local-perspective benchmark of SummEval: scoring two targets separately followed by comparing the scores, and direct pairwise comparison.
@@
Furthermore, we experiment with a more intuitive evaluation approach, where LLMs directly compare each target pair with $I(1)$. As shown in Table 5, it is surprising that the LLMs except those specifically fine-tuned perform worse with direct pairwise comparison than the previous evaluating each target separately and then comparing the scores, which contradicts observations from prior studies (Liusie et al., 2024; Chen et al., 2024). We suppose this is because different evaluation approaches have their respective limits on discriminative power, and differences within the target pairs with $I(1)$ are subtle. Therefore, we further explore target pairs with greater differences up to $I(5)$, as well as other scoring ranges (e.g., 1–100) to observe their influence on the evaluation performance. The results in Figure 3 show that the direct comparison approach is only competitive to the scoring approach when differences between targets are significant. Moreover, LLMs exhibit similar performance when the scoring range extends beyond 1–10, and the optimal range varies.
@@ 6 Related Work
As LLM-as-a-judge methods become increasingly prevalent, prior research has introduced various meta-evaluation benchmarks to assess their effectiveness. Most of these benchmarks emphasize the correlation between LLM-generated evaluations and human judgments (Ye et al., 2024b; Kim et al., 2024a; Zheng et al., 2023; Zeng et al., 2024; Zhang et al., 2023b). Taking a slightly different approach, Thakur et al. (2024) evaluate LLMs using an objective, knowledge-based question-answering task. Tan et al. (2024) further propose a pipeline for automatically constructing benchmarks that focus solely on factual and logical correctness.
In the context of NLG tasks, Bavaresco et al. (2024) explore the performance of various LLM-based evaluators across 20 NLP tasks, including translation and summarization. Other studies have targeted specific NLG tasks, such as translation (Freitag et al., 2024), summarization (Liu et al., 2024c; Siledar et al., 2024), dialogue generation (Zhang et al., 2023a; Mendonça et al., 2024), and story generation (Chhun et al., 2024).
Our work differs from these studies in a fundamental way: we decompose the widely used—yet often ambiguous—measure of human-LLM correlation into two clearer components: the global perspective and local perspective, each with its own dedicated evaluation metric. Additionally, we propose an automatic data construction method to instantiate our meta-evaluation framework. Our approach exposes the limitations of relying solely on human-LLM correlation in NLG meta-evaluation and broader LLM-as-a-judge applications, and introduces a more interpretable alternative.
Beyond human-LLM correlation, Wang et al. (2024a) investigate position bias in pairwise LLM comparisons. Ye et al. (2024a) employ perturbation methods to quantify 12 distinct biases in LLM-based evaluators. Lee et al. (2025) examine the consistency of LLM evaluations under multiple sampling strategies and varying scoring scales. Zhao et al. (2024) explore additional desiderata for pairwise evaluation, such as transitivity. Wang et al. (2024c) design perturbation tests across six NLG tasks to benchmark LLM evaluation capabilities. In contrast to these works, our study remains focused on the correlation between LLMs and human judgments in evaluation tasks.
More closely related to our work are studies in machine translation that seek to redefine the notion of correlation between automatic evaluation metrics and human judgments. However, such efforts are relatively rare and differ significantly in motivation from ours. Specifically, Perrella et al. (2024a) highlight that in machine translation (MT) meta-evaluation, correlation lacks interpretability in terms of range consistency, error attribution, and performance. To address this, they reformulate meta-evaluation as a binary classification and re-ranking task. In contrast, Agrawal et al. (2024) redefine MT meta-evaluation as assessing whether automatic evaluation metrics can effectively distinguish high-quality translations.
@@ 7 Conclusion
Acknowledgements
This work was supported by Beijing Science and Technology Program (Z231100007423011) and Key Laboratory of Science, Technology and Standard in Press Industry (Key Laboratory of Intelligent Press Media Technology). We appreciate the anonymous reviewers for their helpful comments. Xiaojun Wan is the corresponding author.
@@ A.2 Ordinal Classification Metric
Given an ordinal classification task that contains $C$ ordinal classes represented by consecutive integers, let $c_{ij}$ denote the number of items whose gold class is $j$, predicted as $i$ ($i,j\in C$), and $n_{i}$ denote the number of items whose gold class is $i$, and $N$ is the total number of items. Then Closeness Evaluation Measure ($\mathrm{CEM}$) is defined as:
$\mathrm{CEM}=\frac{\sum_{i\in C}\sum_{j\in C}c_{ij}\mathrm{prox}(i,j)}{\sum_{i\in C}n_{i}\mathrm{prox}(i,i)}$
where $\mathrm{prox}(i,j)=-\log\left(\frac{\sum_{k=a}^{b}n_{k}-\frac{n_{i}}{2}}{N}\right)$, with $a=\min(i,j)$ and $b=\max(i,j)$, reflects the informational closeness that the metric assigns to a pair of classes $i,j$. Amigó et al. (2020) conduct extensive analysis and find that $\mathrm{CEM}$ demonstrates good evaluation capability and robustness on ordinal classification, surpassing other common metrics like accuracy and correlation.
@@ A.3 Preliminary Study on SummEval
For our proposed new meta-evaluation framework, we retain only the targets where all three human annotators provide consistent ratings in the original benchmark. For global-perspective meta-evaluation, the results are calculated on an ordinal classification task with five categories and $\mathrm{CEM}$ metric. For local-perspective meta-evaluation, since we lack a fine-grained measure for quality, the target sequence is simply constructed from targets with different original ratings, which actually contradicts our original intent. The corresponding experimental settings are similar to those described in Appendix A.4.
@@ C.1 Construction Process
Since the references in many existing NLG evaluation benchmarks are rule-based rather than human-written, we use the prompts in Table 12 and GPT-4o for generating reference targets for each source to ensure the quality. For global-perspective and local-perspective meta-evaluation, we employ OpenAI o1 to conduct the error injection using the prompts in Tables 18 and 19, respectively. Each error type is randomly selected from the decomposed evaluation sub-aspects in Tables 13 and 14.
@@ C.3 Benchmark Quality
In Table 10, we present the overall comparison results between newly constructed targets and the corresponding targets in the anchor set, based on the estimation method introduced in Section 4.2. More specifically, for the new target $t$ with the rating $r$, columns of $t^{r-1}\prec t$ and $t\prec t^{r+1}$ represent the probability that $t$ has a higher quality than targets $t^{r-1}$ with the rating r1r-1 in the anchor set $A_{r-1}$ and the probability that $t$ has a lower quality than targets $t^{r+1}$ with the rating $r+1$ in the anchor set $A_{r+1}$, respectively. Both these two probabilities are ideally as high as possible. Columns of $t^{r}\prec t$ and $t\prec t^{r}$, on the other hand, indicate the probability that $t$ has a higher and lower quality than targets $t^{r}$ with the same rating $r$ in the anchor set $A_{r}$, respectively, and these two probabilities are expected to be as close as possible. The results demonstrate that our constructed global meta-evaluation benchmarks basically align with our expectations.
Aspect $t^{r-1}\prec t$ $^{r}\prec t tt$$\prec t^{r}$ $t\prec t^{r+1}$
@@ Appendix D Additional Experimental Results
Coherence (Coh): Measure the quality of all sentences of the summary collectively, to fit together and sound naturally. Consider the quality of the summary as a whole. Logical Flow: The sentences in the summary are organized in a logical sequence, ensuring smooth and clear transitions between points of the summary in the given order.
@@
Consistency (Con): Measure whether the facts in the summary are consistent with the facts in the article. Consider whether the summary does reproduce all facts accurately and does not make up untrue information. Factual Accuracy: Each fact stated in the summary accurately reflects the corresponding fact from the news article, without distorted or fabricated information.
@@
Fluency (Flu): Measure the quality of individual sentences of the summary, whether they are well-written and grammatically correct. Consider the quality of individual sentences. Grammatical Correctness: The summary adheres to standard grammar rules without errors in subject-verb agreement, capital letters, tense consistency, or word order.
@@
Relevance (Rel): Measure how well the summary captures the key points of the article. Consider whether all and only the important aspects are contained in the summary. Coverage of Important Information: The summary contains all the key points and information from the news article, without omission (e.g., removing some essential details).
@@
Understandability (Und): Is the response understandable given the previous dialogue context? (Not if it’s on topic, but for example, if it uses pronouns, they should make sense.) Logical Flow: The response is organized in a logical sequence, ensuring smooth and clear transitions between points of the response itself.
@@
Naturalness (Nat): Does the response seem to be something that a person would naturally say? Grammatical Correctness: The response adheres to standard grammar rules without errors in subject-verb agreement, capital letters, tense consistency, word order, or spellings.
@@
Context Maintenance (MCtx): Does the response serve as a valid continuation of the dialogue context (conversation history)? Logical Consistency: The response should logically follow the dialogue context to maintain a smooth continuity and have no contradictions with prior statements or facts in the dialogue context.
@@
Interestingness (Int): Is the response dull or interesting? Content Novelty: The response introduces fresh, unexpected, or unique points and perspectives, which are different from those within the dialogue context, without just repeating the content that the dialogue has mentioned.
@@
Knowledge Use (UK): Given the fact that the response is conditioned on, how well does the response use that fact? Fact Utilization Accuracy: The response should accurately and flawlessly use the information from the given fact, and it must not contain content that conflicts with or distorts and fabricates the given fact.